WARNING: This web is not used anymore. Please use PDBService.MinuteS2August07 instead!
 

Minutes 02 August 07

  • Phone:
    • FNAL: Eric
    • NDGF: Olli
    • PIC: Luis
    • TRIUMF: Denice
    • SARA: Alexander
    • BNL: Carlos
    • ATLAS: Sasha

  • CERN:
    • 3D: Dirk, Eva
    • ATLAS: Gancho, Florbela

  • General announcements:
    • Reminder about the procedures: EGEE broadcast and an entry in the GOC DB (as agreed at the 3D Workshop at SARA) for ALL database interventions.
    • Deployment people is working on a tool to report the interventions.
    • To discuss: transparent intervention should not be announced to everybody

  • Proposal for topics for next mini workshop at PIC:
    • Dates: early-mid October
    • Hands-on activities on adding and removing nodes (see wiki documentation)
    • Security and authentication
    • How to manage the Oracle environment
    • Follow-up from CNAF workshop
    • Database hardware migration
    • Performance and tuning

  • Site status:
    • Tier0 - CERN:
      • Streams problem found when Recyclebin is enabled and drop table operation is replicated (parent-child relation between tables). Workaround: disable the recycle bin – all tier1 sites have already done it. Thanks to all sites for their quick intervention.
      • Investigating previous problem, another one related also with drop table statements have been found. When 2 Streams setup exist between the same source database and same destination database (as it is the case of LFC and LHCb setups), the drop table statement causes an error on the apply side and it is not replicated. Under investigation.
      • Logminer tables maintenance: records from 2 lost old logminer sessions have been removed.
      • CPU July patch incompatible with last Streams patch applied to fix several capture and logminer issues. Merge patch already requested.
      • Regarding the apply problem we had twice at CNAF, apply process getting stuck on apply status, Oracle support does not think this is a Streams problem and they have suggested that it can be some RAC database bugs. I have asked for the corresponding patches.
    • ASGC:
      • Block corruption in one datafile has been fixed. Thanks for everyone help. After fixing this problem, the incremental backup test was successfull.
      • Incremental backup with archivelog everyday configured.
      • CPU July 2007 patch applied to all RAC DB last week. (3D, Castor, FTS2.0 )
      • Installed a new FTS 2.0 server using new rac database. Tests ongoing.
    • RAL:
      • CPU July 2007 patch applied.
      • Disk storage available to be allocated to the existing system.
      • Involved in the BOF sessions at CHEP
    • SARA (Alexander):
      • Nothing to report
      • CPU patch to be applied
    • Triumf (Denice):
      • Nothing to report
      • CPU patch to be applied
    • BNL (Carlos):
      • Nothing to report
      • CPU patch to be applied
      • On holidays next week – mails to general list (see wiki)
    • PIC (Luis):
      • CPU patch to be applied next week: Monday or Tuesday
      • PIC cannot sustained the IO rate load, new hardware, when?
    • NDGF (Olli):
      • Nothing to report
      • CPU patch to be applied next week
    • Proposal for the next security patches: schedule them at Tier1 sites after applying and validating them at Tier0 first. Experiments test patches and potential problems.

  • Experiments status:
    • ATLAS:
      • Rules affecting capture performance, changing the rules order has improved the performance
      • Tests with PVSS, add additional index implies dramatic drop in the performance
      • 5GB of PVSS data replicated per day successfully
      • COOl release out with optimizations. Stress tests to Lyon and CNAF to be repeated
      • Next scalability tests in one month
    • LHCb:
      • No scalability tests yet * Backup and recovery tests pending at CNAF, SARA, Taiwan, Triumf and PIC. RMAN installation is already in place

  • CMS tests:
    • Run tests using calibration. To be checked the load impact on the weekly reports.
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-08-07 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback